OpenStudy (anonymous):

A map of a path is based on a scale of 1:40,000 cm. If the path is 15 cm on the map, how many kilometers is the actual path?

OpenStudy (amistre64):

1 = 40000 --------- 15 = x x = 40000(15) cm

OpenStudy (amistre64):

600,000 6k

OpenStudy (anonymous):

so it's 6k , right?

OpenStudy (amistre64):

yes, id say 6

OpenStudy (anonymous):

okay , can you help me with another one , please?

OpenStudy (amistre64):

maybe, my intelligence waxes and wanes with the phases of the moon lol

OpenStudy (anonymous):

hahaha !

OpenStudy (amistre64):

type faster ;)

OpenStudy (anonymous):

okay, a map of A route has a scale of 1:30,000 cm. IF a lady walks 1.5 km, how many cm is the route on the map?

OpenStudy (amistre64):

1 = 30,000 cm -------- x = 150,000 cm x = 15/3 = 5 cm

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

an ad poster shows a large image of a bottle based on a scale of 7:1 cm. If the actual bottle was 11 cm long, how many cm is it in the poster?

OpenStudy (anonymous):

help? :)

OpenStudy (amistre64):

there you are, i got side tracked lol

OpenStudy (anonymous):

i see.

OpenStudy (amistre64):

7 = 1 ------ x = 11 x = 77 cm

OpenStudy (amistre64):

7(11) = x(1); x = 77

OpenStudy (anonymous):

omg Thank you : )